2. (6 pts) Fill in the missing entries of the matrix, assuming that the equation holds for all values of the variables. ( ) Read Example 5 on page 77 (a)?222-3 2x1-4z2 b) ?? r2-2r3
Answer #1

Solution:

\small 2.

\small \left ( a \right )  

Method I:

Let the required matrix be

adg

Then,

\small \begin{bmatrix} a & b &c \\ d& e &f \\ g & h&i \end{bmatrix}\begin{bmatrix} x_{1}\\ x_{2}\\ x_{3}\\ \end{bmatrix}=\small \begin{bmatrix} 2x_{1}-4x_{2}\\ x_{1}+2x_{2}-x_{3}\\ x_{2}-2x_{3}\\ \end{bmatrix}

210 0 100 c fi adg

\small \therefore\begin{bmatrix} a\\ d\\ g\\ \end{bmatrix}=\small \begin{bmatrix} 2\\ 1\\ 0\\ \end{bmatrix}

In the same way

\small \begin{bmatrix} a & b &c \\ d& e &f \\ g & h&i \end{bmatrix}\begin{bmatrix} 0\\ 1\\ 0\\ \end{bmatrix}=\small \begin{bmatrix} -4\\ 2\\ 1\\ \end{bmatrix}

\small \therefore \begin{bmatrix} b\\ e\\ h\\ \end{bmatrix}=\small \begin{bmatrix} -4\\ 2\\ 1\\ \end{bmatrix}

2 0 001 beh 2 adg

\small \therefore \begin{bmatrix} c\\ f\\ i\\ \end{bmatrix}=\small \begin{bmatrix} 0\\ -1\\ -2\\ \end{bmatrix}

\small \therefore The required matrix is

\small \begin{bmatrix} 2 & -4 &0 \\ 1&2 &-1 \\ 0 &1 & -2 \end{bmatrix}

Method II:

\small \begin{bmatrix} 2x_{1}-4x_{2}\\ x_{1}+2x_{2}-x_{3}\\ x_{2}-2x_{3}\\ \end{bmatrix}=\begin{bmatrix} 2 &-4 & 0\\ 1 & 2 &-1 \\ 0 &1 & -2 \end{bmatrix}\begin{bmatrix} x_{1}\\ x_{2}\\ x_{3}\\ \end{bmatrix}

........................................................................................................................................................................................

\small \left ( b \right )

Method I:

Let the required matrix be

\small \begin{bmatrix} a &b \\ c & d\\ e & f \end{bmatrix}

Then ,

\small \begin{bmatrix} a &b \\ c & d\\ e & f \end{bmatrix}\begin{bmatrix} x_{1}\\ x_{2}\\ \end{bmatrix}=\begin{bmatrix} 2x_{1}-x_{2}\\ x_{1}+4x_{2}\\ x_{2}\\ \end{bmatrix}

\small \begin{bmatrix} a &b \\ c & d\\ e & f \end{bmatrix}\begin{bmatrix} 1\\ 0\\ \end{bmatrix}=\begin{bmatrix} 2\\ 1\\ 0\\ \end{bmatrix}

\small \therefore \begin{bmatrix} a\\ c\\ e\\ \end{bmatrix}=\begin{bmatrix} 2\\ 1\\ 0\\ \end{bmatrix}

\small \begin{bmatrix} a &b \\ c & d\\ e & f \end{bmatrix}\begin{bmatrix} 0\\ 1\\ \end{bmatrix}=\begin{bmatrix} -1\\ 4\\ 1\\ \end{bmatrix}

\small \therefore \begin{bmatrix} b\\ d\\f\\ \end{bmatrix}=\begin{bmatrix} -1\\ 4\\ 1\\ \end{bmatrix}

\small \therefore The required matrix is

\small \begin{bmatrix} 2 &-1 \\ 1& 4\\ 0 & 1 \end{bmatrix}

Method II:

\small \begin{bmatrix} 2x_{1}-x_{2}\\ x_{1}+4x_{2}\\ x_{2}\\ \end{bmatrix}=\begin{bmatrix} 2 &-1\\ 1 & 4\\ 0& 1\end{bmatrix}\begin{bmatrix} x_{1}\\ x_{2}\\ \end{bmatrix}